# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v2

@@ Abstract
Table 1: Layer-resolved $\alpha_{zz}^{(l)}$ in units of $e^{2}/h$ obtained by numerical finite difference and those evaluated by perturbation theory given in the parenthesis.
@@
If $\mathcal{E}_{z}$ is so weak that the electrostatic potential drop between the top and bottom layers can be treated as a perturbation to the system, then $\alpha_{zz}^{(l)}$ can be determined semi-analytically using perturbation theory. After considerable algebra, a semi-analytic formula for $\alpha_{zz}^{(l)}$ is derived. Details are given in Sec. III  of Supplemental Material [31]. We say “semi-analytic” since the determination of the chemical potential still requires solving numerically the system with open boundary condition, which is unavoidable due to the presence of gapless surface states in 3D Chern insulators. Therefore, one can directly calculate $\alpha_{zz}^{(l)}$ using only the energy and Bloch functions of the unperturbed system. As shown in the parenthesis of Table 2, $\alpha_{zz}^{(l)}$ obtained by semi-analytic perturbation theory follow exactly the same quantization rule as those calculated by numerical finite-difference method.
Table 2: Difference between $\alpha_{zz}^{(l)}$ of neighboring layers in units of $e^{2}/h$ obtained by numerical finite difference and those evaluated by perturbation theory given in the parenthesis.
@@
To understand the origin of the quantization in $\alpha_{zz}^{(l)}$, one can consider the simplest case where neighboring layers are completely decoupled from each other such that the Hamiltonian is block diagonal in layer space. The only contribution to $\alpha_{zz}^{(l)}$ comes from $E_{n\textbf{k}}-\mu$ (see Eq. (3)). Then, we can prove exactly that the difference in $\alpha_{zz}^{(l)}$ between the two neighboring layers is quantized (see Sec. V  of Supplemental Material [31])
@@
This quantization rule remains robust even in the presence of considerable interlayer hopping, as shown in Table 2 with $C=1$. Knowing that the total $\alpha_{zz}$ is vanishing, we can deduce that $\alpha_{zz}^{(l)}$ is integer quantized in units of $e^{2}/h$ for odd number of layers or half-integer quantized for even number of layers, consistent with the previous findings.
It is the topological edge states that give rise to the quantized $\partial_{l}\alpha_{zz}(l)$ in a 3D Chern insulator. Intuitively, 3D Chern insulator is formed by stacking $N_{\rm L}$ 2D Chern insulators. There must be $N_{\rm L}$ gapless chiral edge states and a charge neutral $\mu$ crosses with all the edge states. A weak uniform $\mathcal{E}_{z}$ induces different occupation for the edge states of different layer when the system is back to equilibrium. The redistribution of electrons among different layers can be equivalently modelled by layer-resolved chemical potential $\mu^{(l)}=\mu_{0}-e\mathcal{E}_{z}d(l-\Delta l_{N_{\rm L}})$, where $\mu_{0}$ is the chemical potential in the absence of electric field and $\Delta l_{N_{\rm L}}=(N_{\rm L}+1)/2$ such that the total number of electrons is unchanged, namely $\sum\mu^{(l)}=N_{\rm L}\mu_{0}$. Since the chemical potential is in the bulk gap, all the changes in $M_{z}^{(l)}$ stems from the surface-state contribution [6, 30, 32, 33]:
$M_{z}^{(l)}=\mu^{(l)}\frac{e}{hd}C_{z}(l)\,.$ (6)
where the division by $d$ makes $M_{z}^{(l)}$ volumetric. Therefore, we can get a simple formula for $\alpha_{zz}^{(l)}$ as follows
@@
consistent with Eq. (4) knowing that $C_{z}(l)=C$. The behavior for even and odd number of layers are given precisely by the shift $\Delta l_{N_{\rm L}}$. This picture is exact in the interlayer decoupled case and generalizable to the interlayer coupled case by the topological argument, as long as the bulk gap is not closed so that the layer Chern number remains quantized $C_{z}(l)=C$ [34, 35, 36, 37].
Layer-resolved electric polarization–
@@
Before calculating $\alpha_{zz}^{(l)}$, we need to properly define the electronic contribution of layer polarization $P_{z}^{(l)}$, which can be written in terms of the maximally localized Wannier center for each layer in the stacking direction, denoted as $\bar{z}_{l}$ [38]
@@
where $\Omega$ is the area of the 2D magnetic unit cell, interlayer distance is constant $d=1$ and the averaged maximally localized Wannier center is proportional to the averaged Berry phase $\theta_{l}$ calculated by parallel-transport method [38, 39]
$\bar{z}_{l}=\frac{N_{\text{L}}d}{N_{k}}\sum_{k}\frac{-\theta_{l}}{2\pi}\,.$ (9)
where $N_{k}$ is the total number of k points in the $x$-$y$ plane. In practice, parallel-transport method gives $N_{\text{L}}N_{k}$ Berry phases, which are then categorized into $\theta_{l}$ of different layers based on the smallest distance criterion. Here $N_{\rm{L}}$ is the number of layers in an artificially chosen slab, and we consider infinite stacking of slabs with periodic boundary condition imposed, so that the vertical size of 3D Brillouin zone is set to $2\pi/(N_{\rm{L}}d)$.
It turns out that even with a relative strong interlayer coupling ($t_{3},t_{4}$ is of the same order magnitude as $t_{2}$), the Wannier centers still sit closely to each layer [31]. In practice, we use a magnetic field such that the magnetic flux piercing the atomic unit cell is $p/q=1/30$. A $48\times 48\times 48$ k mesh is adopted for numerical calculations. We obtain a similar quantization rule for $\alpha_{zz}^{(l)}$ as shown in Table 3. Compared with previous calculations, the quantization is much better ($\sim 0.1\%$) in such situation with fully periodic boundary condition.
@@
Here we give a somewhat rigorous proof of the quantization rule Eq. (4) for interlayer-coupled 3D Chern insulators. Under finite $B_{z}$, due to Streda formula, the layer charge density $n_{l}$ follows $\partial n_{l}/\partial B_{z}=-C_{z}(l)e^{2}/h$ for all layers, with $C_{z}(l)$ denoting layer Chern number. Therefore, by Streda formula, whenever there is a change in the total charge density (of the $N_{\rm{L}}$-layer unit cell) induced by $B_{z}$, it has to be evenly distributed to the $N_{\rm{L}}$ equidistant averaged Wannier centers, as long as $C_{z}(l)$ is still well defined and quantized. The additional layer charge density leads to a change in layer-resolved polarization $\delta P_{z}^{(l)}=\delta n_{l}\overline{z}_{l}/d$ (according to Eq. (8)), which is induced by $B_{z}$. Thus, the layer-resolved OME coupling is given by:
@@
Note that the average Wannier center $\overline{z}_{l}$ does not necessarily sit at the corresponding layer position due to lack of symmetry. Let $\overline{z}_{l}=ld+\Delta z$ as in our cases with constant interlayer distance $d$, then Eq. (4) follows immediately. Actually the validity of Eq. (4) even does not require uniform layer charge density when $B_{z}=0$. The only requirement is that the bulk gap remains opened and layer Chern number remains exactly quantized as interlayer coupling is turned on, which is obvious for disorder-free 3D Chern insulators where $k_{z}$ is a good quantum number.
For illustration, let us consider a slab breaking inversion and mirror symmetries by introducing a different on-site energy only to the first layer (denoted as substrate layer) such that the total response $\alpha_{zz}$ becomes finite. In practice, the slab consists of five $AA$-stacked layers. We set the on-site energy of the first layer to be $1/6$ while varying the amplitude of $t_{3,4}$. We calculate $\alpha_{zz}^{(l)}$ using Eq. (3) with $90\times 90$ $\mathbf{k}$ mesh. As shown in Table 4, although $\alpha_{zz}^{(l)}$ itself does not follow any quantization rule [31], the difference between the neighboring layer-resolved responses is quantized to layer Chern number as predicted. This shows that the most intrinsic quantization rule is Eq. (4). Nevertheless, symmetry can impose additional constraints on the values of $\Delta z$. For example, mirror or inversion symmetry forces $\Delta z$ to be integer or half integer, as observed in Table 2 and 3.
@@
Therefore, we see from above that the quantization rule Eq. 4 is precisely given by the integer number of transferred states projected to each layer, which essentially originate from the quantized layer Chern number and topological edge states in open boundary condition. This reveals again the topological nature of the quantization in $\alpha_{zz}^{(l)}$.
Robustness of quantization–
@@
In the presence of disorder, as long as the bulk gap is not closed by disorder such that layer Chern number remains well defined and quantized, the change of layer charge density induced by magnetic field would still remain quantized as dictated by Streda formula. Thus, the quantization rule of Eq. (4) is still valid. More numerical evidence of the robustness of Eq. (4) in the presence of disorder is given in Table  5. In practice, we need to calculate $\alpha_{zz}^{(l)}$ numerically by finite difference in the open boundary conditions. Disorder is modelled by random on-site potential. Two types of disorder are introduced. The first type is that on-site energy only fluctuates in-plane repeating itself among all the layers. The second type breaks the translation symmetry such that on-site energy fluctuates among all the sites of all layers. We consider 15 disorder configurations and then analyze the results statistically. Here we set $t_{3}=0.2$, $t_{4}=0$, and a strong disorder with amplitude $0.4t_{3}$. As shown in Table 5, the quantization rule remains good statistically not only for Type-I disorder preserving layer translation symmetry but also Type-II disorder breaking it. This is because the amplitude of the disorder does not close the gap so that layer Chern number can still be well defined.
Discussions–

### v3

@@ II.2 Response of layer-resolved orbital magnetization to vertical electric field
Table 1: Layer-resolved $\alpha_{zz}^{(l)}$ in units of $e^{2}/h$ obtained by numerical finite difference and those evaluated by perturbation theory given in the parenthesis. We consider $AA$,$AB$ stackings and $N_{\rm{L}}=5,6$.
@@
Table 2: Difference between $\alpha_{zz}^{(l)}$ of neighboring layers in units of $e^{2}/h$ obtained by numerical finite difference and those evaluated by perturbation theory given in the parenthesis. We consider $AA$,$AB$ stackings and $N_{\rm{L}}=5,6$.
@@
To understand the origin of the quantization in $\alpha_{zz}^{(l)}$, one can consider the simplest case where neighboring layers are completely decoupled from each other such that the Hamiltonian is block diagonal in layer space. Then, we can prove rigorously that the difference in $\alpha_{zz}^{(l)}$ between the two neighboring layers is exactly quantized (see Sec. V  of Supplemental Materials [29])
@@
This quantization rule remains robust even in the presence of considerable interlayer hopping, as shown in Table 2 with $C=1$. Knowing that the total $\alpha_{zz}$ is vanishing, we can deduce that $\alpha_{zz}^{(l)}$ is integer quantized in units of $e^{2}/h$ for odd number of layers or half-integer quantized for even number of layers, consistent with Table 1.
The quantized $\partial_{l}\alpha_{zz}(l)$ in 3D Chern insulator is attributed to topological edge states. Intuitively, 3D Chern insulator is formed here by stacking $N_{\rm L}$ layers of $C=1$ 2D Chern insulators. There must be $N_{\rm L}$ gapless chiral edge states and a charge neutral $\mu$ crosses with all the edge states. Applying weak uniform $\mathcal{E}_{z}$ induces different occupations for the edge states of different layers when the system is back to equilibrium. The redistribution of electrons among different layers can be equivalently modelled by layer-resolved chemical potential $\mu^{(l)}=\mu_{0}-e\mathcal{E}_{z}d(l-\Delta l_{N_{\rm L}})$, where $\mu_{0}$ is the chemical potential in the absence of electric field and $\Delta l_{N_{\rm L}}=(N_{\rm L}+1)/2$ such that the total number of electrons is unchanged, namely $\sum\mu^{(l)}=N_{\rm L}\mu_{0}$. Since the chemical potential is in the bulk gap, all the changes in $M_{z}^{(l)}$ stems from the surface-state contribution [6, 31, 36, 37, 38]:
$M_{z}^{(l)}=\mu^{(l)}\frac{e}{hd}C_{z}(l)\,,$ (9)
where the division by $d$ makes $M_{z}^{(l)}$ volumetric. Therefore, we can get a simple formula for $\alpha_{zz}^{(l)}$
@@
consistent with Eq. (7) given that $C_{z}(l)=C$. The behavior for even and odd number of layers are precisely given by the shift $\Delta l_{N_{\rm L}}$ as dictated by mirror symmetry. This picture is exact in the interlayer decoupled case and generalizable to the interlayer coupled case by the argument that as long as the bulk gap is not closed, the layer Chern number (see Methods for definition) remains quantized $C_{z}(l)=C$ [34, 39, 33, 40, 35], so that the quantized OME response protected by the layer Chern number should remain robust.
III Layer-resolved electric polarization
@@
Before calculating $\alpha_{zz}^{(l)}$, we need to properly define the electronic contribution of layer polarization $P_{z}^{(l)}$ under weak vertical magnetic field $B_{z}$, which can be written in terms of the maximally localized Wannier center for each layer in the stacking direction, denoted as $\bar{z}_{l}$ [41]
@@
where $\Omega$ is the area of the 2D magnetic primitive cell, interlayer distance $d$ is constant, and the averaged maximally localized Wannier center is proportional to the averaged Berry phase $\phi_{l}$ calculated by parallel-transport method [41, 42]
$\bar{z}_{l}=\frac{N_{\text{L}}d}{N_{k}}\sum_{k}\frac{-\phi_{l}}{2\pi}\,.$ (12)
where $N_{k}$ is the total number of k points in the $x$-$y$ plane. In practice, parallel-transport method gives $N_{\text{L}}N_{k}$ Berry phases, which are then categorized into $\phi_{l}$ of different layers based on the smallest distance criterion [41]. Here $N_{\rm{L}}$ is the number of layers in an artificially chosen repetition slab, namely a supercell in the $z$-direction. We consider infinite stacking of these slabs (formed by 2D Haldane model as before) with periodic boundary condition imposed, so that the vertical size of 3D Brillouin zone is set to $2\pi/(N_{\rm{L}}d)$ with $d$ being interlayer distance.
It turns out that even with a relative strong interlayer coupling (interlayer hopping $t_{3},t_{4}$ is of the same order magnitude as in-plane hopping $t_{2}$), the Wannier centers still sit closely to each layer (see Sec. VI  of Supplemental Materials [29]). In practice, we use a magnetic field such that the magnetic flux piercing the atomic unit cell is $p/q=1/30$. A $48\times 48\times 48$ k mesh is adopted for numerical calculations of the Hofstadter states under magnetic field and the corresponding Berry phases. For the sake of consistency, we also use the same magnetic unit cell in the absence of magnetic field. Then, we can calculate the layer-resolved polarization using Eq. (11). After doing a finite difference between $P_{z}^{(l)}$ in the absence and in presence of magnetic field, we average over the 2D magnetic Brillouin zone to get $\alpha_{zz}^{(l)}$. In the end, we obtain a similar quantization rule for $\alpha_{zz}^{(l)}$ as shown in Table 3 given $C=1$. Comparing with previous calculations, the quantization is much better ($\sim 0.1\%$) in such situation with fully periodic boundary condition, even in the presence of relatively strong interlayer couplings.
@@
Here we prove that the quantization rule Eq. (7) for interlayer-coupled 3D Chern insulators relies on the quantization of layer Chern number. First, due to the Streda formula [43], under finite $B_{z}$, the layer charge density $n_{l}$ follows $\partial n_{l}/\partial B_{z}=-C_{z}(l)e^{2}/h$ for all layers, with $C_{z}(l)$ denoting layer Chern number. Since layer Chern numbers are all equal to each other for a 3D Chern insulator, whenever there is a change in the total charge density (of the $N_{\rm{L}}$-layer unit cell) induced by $B_{z}$, it has to be evenly distributed to the $N_{\rm{L}}$ equidistant averaged Wannier centers, as long as $C_{z}(l)$ is still well defined and quantized. The additional layer charge density leads to a change in layer-resolved polarization $\delta P_{z}^{(l)}=\delta n_{l}\overline{z}_{l}/d$ according to Eq. (11), which is induced by $B_{z}$. Thus, the layer-resolved OME coupling is given by:
@@
Note that the average Wannier center $\overline{z}_{l}$ does not necessarily sit at the corresponding layer position due to lack of crystalline symmetry. However, by translation symmetry in the $z$-direction, we must have $\overline{z}_{l}=ld+\Delta z$ with constant interlayer distance $d$. Then, Eq. (7) follows immediately. Here, it may be more suitable to say “translation symmetry of interlayer coupling”, namely two neighboring layers are always coupled in the same way, because $AB$ stacking does not impose one-layer but two-layer translational symmetry. Nevertheless, the calculations of the Wannier center average over two sublattices so that two stackings make no difference on $\overline{z}_{l}$. So, $\overline{z}_{l}$ must be equidistantly distributed. The importance of translation symmetry can also be seen from Table 2 where the quantization for two outmost layers is worse than the inner layers for $AB$ stacking. Furthermore, the validity of Eq. (7) even does not require uniform layer charge density when $B_{z}=0$. The only requirement is that the bulk gap remains opened and layer Chern number remains exactly quantized as interlayer coupling is turned on, which is obvious for disorder-free 3D Chern insulators where $k_{z}$ is a good quantum number.
IV Robustness of the quantization rule
IV.1 Robustness against symmetry breaking and substrate effects
To illustrate the case without inversion and mirror symmetries, let us go back and consider a slab and introduce a different on-site energy only to the first layer (denoted as “substrate” layer) such that the total response $\alpha_{zz}$ becomes finite due to the breaking of mirror symmetry. For illustration, we consider a slab of five $AA$-stacked layers. We set the on-site energy of the first layer to be $1/6$ while varying the amplitude of $t_{3,4}$. We calculate $\alpha_{zz}^{(l)}$ using Eq. (5) with $90\times 90$ $\mathbf{k}$ mesh. Although $\alpha_{zz}^{(l)}$ itself does not follow any quantization rule (see Table S1 in Supplemental Materials [29]), the difference between the neighboring layer-resolved responses is quantized to layer Chern number as predicted, as shown in Table 4. This shows that the most intrinsic quantization rule is Eq. (7). Nevertheless, symmetry can impose additional constraints on the values of $\Delta z$. For example, mirror or inversion symmetry forces $\Delta z$ to be integer or half integer, which thus enforces $\alpha_{zz}^{(l)}$ to be integer or half integer values, as shown in Table 1 and 3.
@@ IV.1 Robustness against symmetry breaking and substrate effects
Therefore, we see from above that the quantization rule Eq. (7) is precisely given by the integer number of transferred states projected to each layer induced by magnetic field, which essentially originate from the quantized layer Chern number and topological edge states. This reveals again the topological nature of the quantization in $\alpha_{zz}^{(l)}$.
IV.2 Robustness against disorder
@@
In the presence of disorder, as long as the bulk gap is not closed by disorder such that layer Chern number remains well defined and quantized, the change of layer charge density induced by magnetic field would still remain quantized as dictated by the Streda formula. Thus, the quantization rule of Eq. (7) is still valid. More numerical evidence of the robustness of Eq. (7) in the presence of disorder is given in Table  5. In practice, we need to calculate $\alpha_{zz}^{(l)}$ numerically by finite difference in the open boundary conditions. Disorder is modelled by random on-site potential. Two types of disorder are introduced. The first type is that on-site energy only fluctuates in-plane repeating itself among all the layers. The second type breaks the translation symmetry such that on-site energy fluctuates among all the sites of all layers. We consider 15 disorder configurations and then analyze the results statistically. Here we set $t_{3}=0.2$, $t_{4}=0$, and consider a strong disorder with amplitude $0.4t_{3}$. As shown in Table 5, the quantization rule remains good statistically not only for Type-I disorder preserving layer translation symmetry but also Type-II disorder breaking it. This is because the amplitude of the disorder does not close the gap so that layer Chern number can still be well defined.
V Experimental realizations
The considerable and quantized difference between layer-resolved OME $\alpha_{zz}^{(l)}$ is measurable in a slab system. As schematically shown in Fig. 1(c), a slab system can be formed by stacking vertically 2D Chern insulators. The interlayer coupling is in principle tunable, as already experimentally achieved by inserting a spacer with varying thickness between two adjacent Chern layers [44]. Then, the quantized OME response can be experimentally measured in such Chern-insulator slab using state-of-the-art magnetometer techniques such as single-spin scanning magnetometry using nitrogen vacancy in diamond [45, 46] and SQUID-on-tip sensors [47, 48], where “SQUID” is an abbreviation of Superconducting Quantum Interference Device.
V.1 Material proposals
Before providing a specific experimental approach to reveal the topological OME response presented above, we first suggest two possible material systems which can form a slab of quasi-3D Chern insulator. One promising material platform is $\rm{(Bi/Sb)}_{2}\rm{Te}_{3}$ thin film, which is 2D Chern insulator if stacking three Cr-doped quintuple layers [44]. It has been shown in the experiments [44] that quasi-3D Chern insulator slab can be synthesized by alternately growing three Cr-doped quintuple layers and four undoped ones (served as spacer) in the $z$-direction, which exhibit well quantized anomalous Hall resistance with high Chern numbers thanks to good quality of samples.
Another candidate is the Mn(Bi/Sb)2Te4[(Bi/Sb)2Te3]n-class ($n=0,1,2,...$) of materials. To get a building block of 3D Chern insulator, we explore several kinds of Mn(Bi/Sb)2Te4[(Bi/Sb)2Te3]n-type thin films with different number of layers and different in-plane strain magnitudes using first principles DFT calculations (see Sec. VIII of Supplemental Materials). As shown in Table 6, the most stable case is 3-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ in ferromagnetic state, which is a $C=1$ Chern insulator and remains topologically nontrivial over a large range of in-plane strain amplitudes from -2% to 2%. The indirect gap of such Chern insulator is as large as 464646\,46meV as shown in Fig. 2(a). Such a large topological gap makes 3-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ a promising building block for quasi-3D Chern-insulator slab. For illustration, we construct the Wannier tight-binding Hamiltonian (based on DFT calculations) for 3, 6, and 9-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ slabs assuming a proper spacer is inserted between every two adjacent three-layer $C=1$ entities, so that the inter-entity coupling can be determined from DFT calculations (see Sec. VIII of Supplemental Materials). As expected, the calculated total Chern numbers for the 3, 6, 9-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ slabs are $C=1,2,3$, respectively. The calculated layer Chern numbers are always well quantized to 1 as shown in Fig. 2(b), even under substantial in-plane strain. This suggests that MnBi2Te4 thin films with properly intercalated interlayer spacer is an ideal platform to test the quantization rule Eq. (7).
Table 6: Chern numbers of Mn(Bi/Sb)2Te4[(Bi/Sb)2Te3]n-class thin films under different strain magnitudes. “FM“ and “AFM” stand for interlayer ferromagnetic and anti-ferromagnetic configurations, respectively.
$\mathrm{system\,(layers\mid magnetism\mid strain\,axis)}$2.0% -1.5% -1.0% -0.5% 0.0% 0.5% 1.0% 1.5% 2.0%
[…]
Figure 2: (a) Band structure of 3L MBT without strain in which the zero energy indicates chemical potential and the indirect gap is 46 meV. (b) Layer Chern number and total Chern number for 3,6,9-layer MBT in the presence of in-plane strain -2%, 0% and 2%. The height of each sub-bar indicates the calculated layer Chern number for each 3L MBT entity using Eq. (6) under the same $150\times 150$ $\mathbf{k}$-mesh. The total height is then the total Chern number. The error of the quantization is given above each bar. We see that the quantization of the calculated Chern numbers is better with strain 2% but worse with strain -2% because the indirect gap increases with strain from -2% to 2% (see Sec. VIII  of Supplemental Materials [29]).
V.2 Workflow for experimental detection
Once the encapsulated device of quasi-3D Chern insulator slab is synthesized, one can use a magnetometer with high precision to acquire profiles of magnetic field $\mathbf{B}$ outside the device. The sensitivity of the state-of-art technique such as single-spin scanning magnetometry using nitrogen vacancy in diamond [45, 46] and SQUID-on-tip sensors [47, 48] is below nano Tesla [49], through which topological magnetoelectric effect in the quantum Hall state has already been observed [47]. Then, layer-resolved magnetization $M_{z}^{(l)}$ can be reconstructed by reverse propagation [50, 51, 30, 52]. We briefly discuss how this method can be implemented for our problem.
We suppose that the measured magnetic field is solely generated by the magnetization of the device. The law of Biot-Savart in magnetostatics stipulates that
$\mathbf{B}=\frac{\mu_{0}}{4\pi}\iiint d^{3}r^{\prime}\,\frac{\mathbf{J}(\mathbf{r}^{\prime})\times(\mathbf{r}-\mathbf{r}^{\prime})}{|\mathbf{r}-\mathbf{r}^{\prime}|^{3}}$ (14)
with $\mu_{0}$ is the permeability of vacuum and the current density $\mathbf{J}(\mathbf{r}^{\prime})$ is bound current generated by magnetization $M$, $\mathbf{J}=\nabla\times\mathbf{M}$. Here, we neglect in-plane magnetization since Chern layers are atomically thin, and out-of-plane magnetization contributed by each Chern layer depends only on spatial coordinates $x$ and $y$, $\mathbf{M}=M(x,y)\hat{z}$. Let us take SQUID-on-tip technique as example, which probes only the $z$ component of magnetic field. Direct calculations from Eq. (14) lead to a formula for magnetic field outside the device created by a Chern layer confined between $z_{1}$ and $z_{2}$
$B_{z}(x,y,z)\;$
[…]
with the kernel function $K(x,y,z)$
$K(x,y,z)\;$
$=$ $\frac{\mu_{0}}{4\pi}\left(\frac{z-z_{2}}{(x^{2}+y^{2}+(z-z_{2})^{2})^{\frac{3}{2}}}-\frac{z-z_{1}}{(x^{2}+y^{2}+(z-z_{1})^{2})^{\frac{3}{2}}}\right).$ (16)
Eq. (V.2) can be nicely dealt under Fourier transformation such that convolution becomes trivial product
$\tilde{B}_{z}(k_{x},k_{y},z)=\tilde{K}(k_{x},k_{y},z)\,\tilde{M}(k_{x},k_{y})$ (17)
where the notation with tilde is the Fourier transform of the corresponding function. It turns out that the kernel function $\tilde{K}(k_{x},k_{y},z)$ can be computed analytically
$\tilde{K}(k_{x},k_{y},z)=\frac{\mu_{0}}{2}e^{-kz}\left(e^{kz_{2}}-e^{kz_{1}}\right)$ (18)
with $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$. Therefore, as long as the profile of magnetic field at given $z$ is known, we can calculate its Fourier transform and then invert the kernel function in Eq. (17) to get the profile of the magnetization in the Chern layer after inverse Fourier transformation. This process is called reverse propagation.
Nevertheless, the measured magnetic field $B_{z}$ is generated by all the Chern layers in quasi-3D Chern insulator slab. Each Chern layer $l$ is equipped with its own kernel function $\tilde{K}_{l}$ where $z_{1,2}$ are replaced by the layer-resolved $z_{1,2}^{(l)}$ in Eq. (18). So, Eq. (17) becomes
$\tilde{B}_{z}(k_{x},k_{y},z)=\sum_{l=1}^{N_{\text{L}}}\tilde{K}_{l}(k_{x},k_{y},z)\,\tilde{M}_{z}^{(l)}(k_{x},k_{y})$ (19)
where $N_{\text{L}}$ is the total number of Chern layers and $\tilde{M}_{z}^{(l)}$ is the Fourier transform of magnetization of layer $l$. To get $\tilde{M}_{z}^{(l)}$, we need to collect the data of $B_{z}$ at $N_{\text{L}}$ different height $h_{i}$ with $i=1,\dots,N_{\text{L}}$ outside the device. Then, reverse propagation becomes matrix inversion
$\tilde{M}_{z}^{(l)}(k_{x},k_{y})$ $=\sum_{i=1}^{N_{l}}[\Gamma^{-1}(k_{x},k_{y})]_{l,i}\,\tilde{B}_{z}(k_{x},k_{y},h_{i})$ (20)
with $\Gamma_{i,l}(k_{x},k_{y})=\tilde{K}_{l}(k_{x},k_{y},h_{i})$. Finally, using reverse propagation method, we can record the change in layer-resolved magnetization in 3D Chern insulator slabs after applying a non-disruptive out-of-plane electric field via gating. The quantization rule Eq. (7) of OME coupling can be then measured by finite difference.
VI Discussions
In real materials, each Chern layer has a finite thickness, so that we need to replace the interlayer distance by the lattice constant in the $z$ direction in all the results above. In particular, our proof by the Streda formula of the quantization rule Eq. (7) is still valid since it relies only on the translation symmetry and the quantization of layer Chern number. We also note that Seleznev et al. [38] have recently proposed a new definition for the local marker of orbital magnetization, leading to another definition of layer-resolved orbital magnetization that differs from the one adopted in this work. Nevertheless, we have checked that our proposed quantization rule remains valid under this alternative definition of layer-projected orbital magnetization (see Sec. IX in Supplemental Materials).
In fact, by adopting the “gauge-discontinuity” formalism of Chern-Simons OME coupling [28], the quantization rule introduced in this work can be derived in a mathematically rigorous way even in the presence of considerable interlayer coupling. It can be interpreted as a new type of “anomalous” Chern-Simons OME coupling unique to 3D Chern insulators. We refer the readers to Ref. 53 for the full mathematical derivations.